# The unknotting number, hard unknot diagrams, and reinforcement learning

### v1

@@ 1. Introduction
Knot theory plays a fundamental role in low-dimensional topology. A knot is a smooth embedding $K\colon S^{1}\hookrightarrow S^{3}$. We say that the knots $K$ and $K^{\prime}$ are equivalent if there is an orientation-preserving automorphism $\varphi$ of $S^{3}$ such that $\varphi\circ K=K^{\prime}$. We can represent a knot using a projection onto $S^{2}$ with only transverse double point singularities, together with information at each double point about which strand is higher. This is called a knot diagram. Two knot diagrams represent the same knot if and only if they are related by a sequence of Reidemeister moves R1–R3. For textbooks on knot theory, see Burde–Zieschang [6], Lickorish [23], and Rolfsen [33].
@@
A more intrinsic definition of the unknotting number is obtained using crossing arcs. A crossing arc $a$ for a knot $K$ is a framed, oriented arc smoothly embedded in $S^{3}$ such that $K\cap a=\partial a$. A crossing change along $a$ is obtained by performing a finger move on $K$ along $a$. Then $u(K)$ is the minimal number of such crossing changes that result in $U$. If $a_{1},\dots,a_{n}$ is a collection of crossing arcs, then we can always isotope them and the knot $K$ so that the crossing arcs become short, vertical line segments, in which case the crossing changes along them correspond to crossing changes in the respective diagram. This shows the equivalence between the two definitions. Yet another description of the unknotting number is given by taking the minimal number of double points that appear in a generic regular homotopy from $K$ to $U$; see Lickorish [23, p. 7].
@@
If $c(\mathcal{D})=c(K)$, we say that $\mathcal{D}$ is a minimal crossing number diagram of $K$. Note that a knot can have several minimal crossing number diagrams. There are finitely many knots for each crossing number, so knots are usually tabulated by crossing number. See Rolfsen [33] for a table of knots up to ten crossings and KnotInfo [24] for knots up to 13 crossings. For example, the only crossing number zero knot is the unknot $U$ and there are no knots of crossing number one or two. There are two knots of crossing number three, the right-handed trefoil $3_{1}$ (often denoted by $T_{2,3}$, as it is the $(2,3)$-torus knot) and its mirror, the left-handed trefoil $-3_{1}$. There is only one knot of crossing number four, the figure eight knot $4_{1}$, which is equivalent to its mirror. The number of knots grows exponentially as the crossing number increases. For an integer $c\in[3,10]$, the notation $c_{n}$ refers to knot number $n$ of crossing number $c$ in Rolfsen’s table.
@@
A conjecture of Bernhard [3] and Jablan [17] stated that every knot $K$ has a minimal crossing number diagram $\mathcal{D}$ and a crossing $c$ such that changing $c$ results in a knot $K^{\prime}$ with $u(K^{\prime})=u(K)-1$. If true, this would yield an algorithm for computing $u(K)$. However, Brittenham and Hermiller [5] have shown this to be false: at least one of 12n288, 12n491, 12n501, and 13n3370 violates the conjecture. One can obtain each of 12n288, 12n491, and 12n501 from 13n3370 via a single crossing change. The knot 13n3370 is the closure of a 20-crossing braid, where changing a single crossing gives 11n21 that has unknotting number one. So $u(13n3370)\leq 2$, but it is hard to find a diagram $\mathcal{D}$ of 13n3370 with $u(\mathcal{D})=2$ by applying random Reidemeister moves to a minimal crossing number diagram.
@@
In practice, one can often get a good upper bound on $u(K)$ by simplifying the diagram, changing a crossing such that the crossing number is minimal after simplifying, and repeating this process.
@@
Most known lower bounds on $u(K)$ are also lower bounds on $g_{4}(K)$. Among these, $|\sigma(K)|/2$, $|\tau(K)|$, $|\nu(\pm K)|$, and $|s(K)|/2$ are efficiently computable, where $\sigma$ is the signature [18, Proposition 4.28], the invariants $\tau$ and $\nu$ were defined and shown to bound $g_{4}(K)$ by Ozsváth and Szabó [28] using knot Floer homology [29][32], and $s$ was defined and shown to bound $g_{4}(K)$ by Rasmussen [31] via Khovanov homology [19]. We obtain $u(K)$ if the upper and lower bounds agree. For example, Kronheimer and Mrowka [20] have shown that, for the torus knot $T_{p,q}$, we have
@@ 1.2. Additivity of unknotting number
To find counterexamples to the conjecture, one needs to start with knots $K$ and $K^{\prime}$ with known unknotting numbers, and then to find efficient ways of unknotting $K\#K^{\prime}$. One significant source of knots $K$ with known unknotting number is those for which $u(K)=|\sigma(K)|/2$. Given two such knots $K$ and $K^{\prime}$, then of course $u(K\#K^{\prime})=u(K)+u(K^{\prime})$ if $\sigma(K)$ and $\sigma(K^{\prime})$ have the same signs. However, if they have opposite signs, then there is no obvious reason why $K\#K^{\prime}$ cannot be a counterexample to the conjecture. A further source of knots with known unknotting number are torus knots, and again there seems to be no known reason why $u(K\#K^{\prime})=u(K)+u(K^{\prime})$ for torus knots of signature with opposite signs. Indeed, it is currently unknown whether
@@ 1.3. Finding efficient unknotting sequences
A crucial part of the strategy for disproving Conjecture 1.1 is to be able to find short unknotting sequences. In particular, in the case of $K\#K^{\prime}$, the number of crossing changes needs to be less than $u(K)+u(K^{\prime})$. Even when one is presented with a diagram $\mathcal{D}$ for a knot $K$, it is not straightforward to compute $u(\mathcal{D})$ when the crossing number of $\mathcal{D}$ is large. For a knot diagram $\mathcal{D}$ with $n$ crossings, $u(\mathcal{D})\leq n/2$, hence there are at least $2^{n-1}$ possibilities for the subset of crossings that yield a diagram of the unknot. This makes computing $u(\mathcal{D})$ practically impossible when $n$ is large.
@@
By combining the agent with lower bounds coming from invariants such as the signature, $\tau$, $\nu$, and $s$, we have obtained a dataset of about 57k knot diagrams with known unknotting numbers. We have then taken connected sums of such diagrams, which were overlaid and, in some cases, then randomly mixed using Reidemeister moves. We have also run it on connected sums of braid closures that were mixed by inserting subwords representing the trivial braid. The agent found unknotting sequences that involved several crossing changes that resulted in hyperbolic knots, and were hence not connected sums. This has led us to diagrams of connected sums of knots $K$ and $K^{\prime}$ that admit an unknotting subset of crossings of size $u(K)+u(K^{\prime})$, such that any single crossing change from the unknotting subset results in a hyperbolic knot, hence disproving Conjecture 1.2. In fact, we will prove the following:
@@ 2. Some background on Machine Learning
In SL, we are given a labelled dataset. In other words, we know the values of a function at certain points. We split our dataset into a training set, which is typically about 80%, and a test set. We would like to learn, or approximate the function only using the training set such that the error (e.g., $L^{2}$-norm) is small on the whole dataset.
The most classical example is linear regression. More generally, Hornik, Stinchcombe, and White [15] have shown that neural networks (NNs) are universal function approximators, if one is allowed to vary the architecture. A neural network is a composition of a sequence of affine maps and some simple non-linearities in between, such as $\max(0,x)$ applied coordinate-wise. The network is trained using some variant of stochastic gradient descent. One initialises the affine maps, for example, randomly, then computes an approximate of the gradient of the error on a subset of the training set (whose cardinality is called the batch size), and changes the affine maps in the direction of the gradient according to some step size (or learning rate). This is repeated a number of times, and a pass through the whole training set is called an epoch.
There have been several of applications of SL to knot theory in recent years, mostly aimed at finding connections between knot invariants. See, for example, Hughes [16] and Davies et al. [9].
RL is a machine learning paradigm where an agent (in our case, a computer software) learns to perform actions to maximise a cumulative reward while interacting with some environment. Typical examples are provided by the games of chess and Go, self-driving cars, and humanoid robots that learn to walk. Training a SL model is often much simpler than RL. There have been only two applications of RL to topology so far. Gukov, Halverson, Ruehle, and Sułkowski [13] focused on unknot recognition. Furthermore, Gukov, Halverson, Manolescu, and Ruehle [12] have developed RL agents that search for ribbon disks for a knot. In this rest of this section, we give an overview of RL and imitation learning.
@@
Mathematically, RL can be phrased as a Markov decision process, which is a tuple $(S,A,P_{a},R_{a})$, where
@@
In our case, $S$ consists of certain invariants of diagrams that can be obtained by crossing changes from a fixed knot diagram $\mathcal{D}$. An action is changing a crossing of $\mathcal{D}$. A crossing change is deterministic, so $P_{a}(s,s^{\prime})$ is 1 if a crossing change $a\in A_{s}$ results in $s^{\prime}$ and is 0 otherwise. If $a\in A_{s}$ leads to $s^{\prime}$, then the reward $R_{a}(s,s^{\prime})$ is $1$ if $s^{\prime}$ is a diagram of the unknot $U$ and is 00 otherwise.
@@
A classical approach to solving Markov decision process is Q-learning, where ‘Q’ stands for ‘quality’. Its goal is to learn the state-action value $Q(s,a)$, which is the expected discounted total reward if action $a\in A_{s}$ is taken in state $s\in S$. At time $t$, the agent selects action $a_{t}$, observes a reward $r_{t}$, and enters state $s_{t+1}$. We initialise $Q$ randomly and updated it via the Bellman equation
@@
for the basis vector $e_{s}$ of $\mathbb{R}^{S}$ corresponding to the state $s\in S$ and the basis vector $e_{a}$ of $\mathbb{R}^{A}$ corresponding to the action $a\in A$. The weights of the network are updated using the Bellman equation.
@@ 3.1. Imitation learning and unknotting
We used behavioural cloning based on a NN to predict for each crossing the probability that it lies in a minimal unknotting set. If the predicted probability for a crossing $c$ is larger than $0.5$, then we interpret this such that $c$ does lie in a minimal unknotting set. Expert data was obtained from brute-forced minimal unknotting sets of knots up to 30 crossings.
See Figure 1 for the percentage of the training set for which a given crossing appears in a minimal unknotting set. The distribution of the percentage of crossings that lie in a minimal unknotting set for diagrams in the dataset is shown in Figure 2. The unknotting numbers of the diagrams ranges between 1 and 8.
Figure 1. For each crossing, the percentage of the training set for which it appears in a minimal unknotting set.
@@
To find out which subset of invariants to use, we set up a Supervised Learning experiment to predict whether a crossing lies in a minimal unknotting sequence. We considered neural networks with hidden layers of size [256, 256], [1024, 1024, 1024], and [2048, 2048, 2048], respectively. Learning was done in 10k steps, with learning rate 0.01 or 0.005, momentum 0.9, and batch size 2048. Accuracy ranged between 84.6% and 88.1%, with baseline 50.1% when predicting the most commonly occurring crossing in unknotting sets; see Figure 1 (crossing 2 is in a minimal unknotting set for 50.1% of diagrams in the dataset). Highest accuracy was achieved by hidden layers of size [1024,1024,1024] and learning rate 0.01 using all features. When only using the features in (3) on the same architecture, we obtained almost identical, 87.97% accuracy.
@@ 3.2. Reinforcement learning and unknotting
Initially, we observed that the agent frequently revisited the same diagram. If a crossing appeared multiple times in an unknotting sequence, we counted it modulo 2. Furthermore, we experimented with disallowing revisiting the same Alexander or Jones polynomial, which decreased the percentage of unsolved diagrams for 50–100-crossing knots from 0.49% to 0.05%. There was also a small increase in performance when disallowing Jones revisits. When using an agent based on Alexander polynomial features only, disallowing Alexander polynomial revisits substantially improved performance, except when solving connected sums with forced inter-component crossings. When we allowed revisits, the agent sometimes found shorter unknotting sequences (after counting the number of times each crossing change was made modulo 2). The Alexander only policy was 10 times faster than the one using the Jones polynomial. Running the agent 10 times on each diagram substantially increased performance.
@@
We also compared the Jones unknotting agent to various baselines. A naive baseline of randomly switching crossings is fast and simple to implement, but unlikely to find minimal unknotting sequences for diagrams with many crossings. Similarly to [12], we developed a Bayesian Optimisation (BO) based random agent. The agent chooses randomly from the following action categories: Reidemeister I move, reverse Reidemeister I move, Reidemeister II move, reverse Reidemeister II move, Reidemeister III move, and crossing switch. The random agent chooses one of these action categories by randomly sampling according to certain weights. The concrete action is then sampled uniformly from all valid moves from that category, e.g. from all valid crossing indices. We optimised the sampling weights using BO using the average reward on a heldout validation set as the fitness metric. We limited the weights to be integers between $1$ and $100$. The best fitness was achieved by weight vectors where weights for simplifying moves (Reidemeister I or II) as well as neutral moves (Reidemeister III) were near $100$, while weights for moves which increase the crossing count (reverse Reidemeister I and II) or which incur a reward penalty (crossing switch) were assigned a weight near $1$. Intuitively, this means the optimisation procedure converged to a solution which always attempts to maximally simplify the knot diagram before making a crossing switch. We then implemented further baselines which follow this insight of first simplifying then switching a crossing. Simplify (random) utilizes SnapPy to first simplify the knot diagram before a random crossing is switched. Simplify (min. crossing) always switches the crossing after which the resulting diagram has the least crossings after simplification. Figure 4 shows how different unknotting strategies compare on a test set of $100,000$ diagrams. The RL agent utilizing the Jones polynomial manages to unknot more knots than the other strategies and is also capable of producing more efficient unknotting sequences.

### v2

@@ 1. Introduction
Knot theory plays a fundamental role in low-dimensional topology. A knot is a smooth embedding $K\colon S^{1}\hookrightarrow S^{3}$. We say that the knots $K$ and $K^{\prime}$ are equivalent if there is an orientation-preserving automorphism $\varphi$ of $S^{3}$ such that $\varphi\circ K=K^{\prime}$. We can represent a knot using a projection onto $S^{2}$ with only transverse double point singularities, together with information at each double point about which strand is higher with respect to the projection. This is called a knot diagram. Two knot diagrams represent the same knot if and only if they are related by a sequence of Reidemeister moves R1–R3. For textbooks on knot theory, see Burde–Zieschang [6], Lickorish [23], and Rolfsen [33].
@@
A more intrinsic definition of the unknotting number is obtained using crossing arcs. A crossing arc $a$ for a knot $K$ is a framed, oriented arc smoothly embedded in $S^{3}$ such that $K\cap a=\partial a$. Here, a framing of $a$ is a trivialisation $(v_{1},v_{2})$ of the normal bundle of $a$ such that $v_{1}(a(0))$ and $v_{2}(a(1))$ are positive tangents to $K$. A crossing change along $a$ is obtained by performing a finger move on $K$ along $a$. More precisely, for some $\varepsilon>0$ small, we replace a small segment of $K$ passing through $a(0)$ with curves $\varepsilon v_{1}(a(t))$ and $-\varepsilon v_{1}(a(t))$ for $t\in I$, together with a semi-circular arc connecting $\varepsilon v_{1}(a(1))$ and $-\varepsilon v_{1}(a(1))$. Then $u(K)$ is the minimal number of such crossing changes that result in $U$. If $a_{1},\dots,a_{n}$ is a collection of crossing arcs, then we can always isotope them and the knot $K$ so that the crossing arcs become short, vertical line segments with parallel framing, in which case the crossing changes along them correspond to crossing changes in the respective diagram. This shows the equivalence between the two definitions. Yet another description of the unknotting number is given by taking the minimal number of double points that appear in a generic regular homotopy from $K$ to $U$; see Lickorish [23, p. 7].
@@
If $c(\mathcal{D})=c(K)$, we say that $\mathcal{D}$ is a minimal crossing number diagram of $K$. Note that a knot can have several inequivalent minimal crossing number diagrams. There are finitely many knots for each crossing number, so knots are usually tabulated by crossing number. See Rolfsen [33] for a table of knots up to ten crossings and KnotInfo [24] for knots up to 13 crossings. For example, the only crossing number zero knot is the unknot $U$ and there are no knots of crossing number one or two. There are two knots of crossing number three, the right-handed trefoil $3_{1}$ (often denoted by $T_{2,3}$, as it is the $(2,3)$-torus knot) and its mirror, the left-handed trefoil $-3_{1}$. There is only one knot of crossing number four, the figure eight knot $4_{1}$, which is equivalent to its mirror. The number of knots grows exponentially as the crossing number increases; see Welsh [38]. For an integer $c\in[3,10]$, the notation $c_{n}$ refers to knot number $n$ of crossing number $c$ in Rolfsen’s table.
@@
A conjecture of Bernhard [3] and Jablan [17] stated that every knot $K$ has a minimal crossing number diagram $\mathcal{D}$ and a crossing $c$ such that changing $c$ results in a knot $K^{\prime}$ with $u(K^{\prime})=u(K)-1$. If true, this would yield an algorithm for computing $u(K)$: Enumerate all minimal crossing number diagrams of $K$, list all diagrams that can be obtained from these by changing one crossing, and repeat with these knots until we first reach the unknot. The number of steps is $u(K)$. However, Brittenham and Hermiller [5] have shown this to be false: at least one of 12n288, 12n491, 12n501, and 13n3370 violates the conjecture. One can obtain each of 12n288, 12n491, and 12n501 from 13n3370 via a single crossing change. The knot 13n3370 is the closure of a 20-crossing braid, where changing a single crossing gives 11n21 that has unknotting number one. So $u(13n3370)\leq 2$, but it is hard to find a diagram $\mathcal{D}$ of 13n3370 with $u(\mathcal{D})=2$ by applying random Reidemeister moves to a minimal crossing number diagram.
@@
In practice, a good upper bound on $u(K)$ can often be obtained by first simplifying a diagram of $K$ (removing crossings using Reidemeister moves) and then selecting a crossing to change by picking one that results in a diagram that can be further simplified to have the fewest crossings. This process is then repeated iteratively.
@@
Most known lower bounds on $u(K)$ are also lower bounds on $g_{4}(K)$. Among these, $|\sigma(K)|/2$, $|\tau(K)|$, $|\nu(\pm K)|$ (where $-K$ is the mirror of $K$), and $|s(K)|/2$ are efficiently computable, where $\sigma$ is the signature [18, Proposition 4.28], the invariants $\tau$ and $\nu$ were defined and shown to bound $g_{4}(K)$ by Ozsváth and Szabó [28] using knot Floer homology [29][32], and $s$ was defined and shown to bound $g_{4}(K)$ by Rasmussen [31] via Khovanov homology [19]. (Note that $\sigma$, $\tau$, and $s$ are symmetric with respect to mirroring, but $\nu$ is not.) We obtain $u(K)$ if the upper and lower bounds agree. For example, Kronheimer and Mrowka [20] have shown that, for the torus knot $T_{p,q}$, we have
@@ 1.2. Additivity of unknotting number
To find counterexamples to the conjecture, one needs to start with knots $K$ and $K^{\prime}$ with known unknotting numbers, and then to find efficient ways of unknotting $K\#K^{\prime}$. One significant source of knots $K$ with known unknotting number is those for which $u(K)=|\sigma(K)|/2$. Given two such knots $K$ and $K^{\prime}$, then of course $u(K\#K^{\prime})=u(K)+u(K^{\prime})$ if $\sigma(K)$ and $\sigma(K^{\prime})$ have the same signs, as $\sigma$ is additive under connected sum. However, if they have opposite signs, then there is no obvious reason why $K\#K^{\prime}$ cannot be a counterexample to the conjecture. A further source of knots with known unknotting number are torus knots, and again there seems to be no known reason why $u(K\#K^{\prime})=u(K)+u(K^{\prime})$ for torus knots of signature with opposite signs. Indeed, it is currently unknown whether
@@ 1.3. Finding efficient unknotting sequences
A crucial part of the strategy for disproving Conjecture 1.1 is to be able to find short unknotting sequences. In particular, in the case of $K\#K^{\prime}$, the number of crossing changes needs to be less than $u(K)+u(K^{\prime})$. Even when one is presented with a diagram $\mathcal{D}$ for a knot $K$, it is not straightforward to compute $u(\mathcal{D})$ when the crossing number of $\mathcal{D}$ is large. For a knot diagram $\mathcal{D}$ with $n$ crossings, $u(\mathcal{D})\leq n/2$. If a set of crossings is unknotting, so is its complement. Hence, there are $2^{n-1}$ possibilities for the subset of crossings that yield a diagram of the unknot, up to taking complements. This makes computing $u(\mathcal{D})$ practically impossible when $n$ is large.
@@
By combining the agent with lower bounds coming from invariants such as the signature, $\tau$, $\nu$, and $s$, we have obtained a dataset of about 57k knot diagrams with known unknotting numbers. We have then taken connected sums of such diagrams, which were overlaid and, in some cases, then randomly mixed using Reidemeister moves. We have also run it on connected sums of braid closures that were mixed by inserting subwords representing the trivial braid. The agent found unknotting sequences that involved several crossing changes that resulted in hyperbolic knots, and were hence not connected sums. Throughout this paper, we will use the terminology that a crossing change in a diagram of a connected sum is inter-component if it results in a knot that is not a connected sum (e.g., hyperbolic) and is in-component otherwise. This led us to diagrams of connected sums of knots $K$ and $K^{\prime}$ that admit an unknotting subset of crossings of size $u(K)+u(K^{\prime})$, such that any single crossing change from the unknotting subset results in a hyperbolic knot, hence disproving Conjecture 1.2. In fact, we will prove the following:
@@ 2. Some background on Machine Learning
In SL, we are given a labelled dataset. In other words, we know the values of a function at certain points. We split our dataset into a training set, which is typically about 80%, and a test set. We would like to learn, or approximate, the function using only the training set such that the error (e.g., $L^{2}$-norm) is small on the whole dataset.
The most classical example is linear regression. More generally, Hornik, Stinchcombe, and White [15] have shown that neural networks (NNs) are universal function approximators, if one is allowed to vary the architecture. A neural network is a composition of a sequence of affine maps and some simple non-linearities in between, such as $\max(0,x)$ applied coordinate-wise. The network is trained using some variant of stochastic gradient descent. One initialises the affine maps, for example, randomly, then computes an approximation of the gradient of the error on a subset of the training set (whose cardinality is called the batch size), and changes the affine maps in the direction of the negative gradient according to some step size (or learning rate). This is repeated a number of times, and a pass through the whole training set is called an epoch.
There have been several applications of SL to knot theory in recent years, mostly aimed at finding connections between knot invariants. See, for example, Hughes [16] and Davies et al. [9].
RL is a machine learning paradigm where an agent (in our case, computer software) learns to perform actions to maximise a cumulative reward while interacting with an environment. Typical examples are provided by the games of chess and Go, self-driving cars, and humanoid robots that learn to walk. Training a SL model is often much simpler than RL. There have been only two applications of RL to topology so far. Gukov, Halverson, Ruehle, and Sułkowski [13] focused on unknot recognition. Furthermore, Gukov, Halverson, Manolescu, and Ruehle [12] have developed RL agents that search for ribbon disks for a knot. In this rest of this section, we give an overview of RL and imitation learning.
@@
Mathematically, RL is a framework which can be used to find solutions to problems that are formulated as Markov decision processes. A Markov decision process is a tuple $(S,A_{s},P_{a},R_{a})$, where
@@
In our case, $S$ consists of certain invariants of diagrams that can be obtained by crossing changes from a fixed knot diagram $\mathcal{D}$. An action is changing a crossing of $\mathcal{D}$. A crossing change is deterministic, so $P_{a}(s,s^{\prime})$ is 1 if a crossing change $a\in A_{s}$ is applied to the diagram associated to state $s$ and results in the state $s^{\prime}$ and is 0 otherwise. If $a\in A_{s}$ leads to $s^{\prime}$, then the reward $R_{a}(s,s^{\prime})$ is $1$ if $s^{\prime}$ is a diagram of the unknot $U$ and is 00 otherwise.
@@
A classical approach to solving Markov decision processes is Q-learning, where ‘Q’ stands for ‘quality’. Its goal is to learn the state-action value $Q(s,a)$, which is the expected discounted total reward if action $a\in A_{s}$ is taken from state $s\in S$, and the policy is followed thereafter. At time $t$, the agent selects action $a_{t}$, observes a reward $r_{t}$, and enters state $s_{t+1}$. We initialise $Q$ randomly and updated it via the Bellman equation
@@
for the basis vector $e_{s}$ of $\mathbb{R}^{S}$ corresponding to the state $s\in S$ and the basis vector $e_{a}$ of $\mathbb{R}^{A}$ corresponding to the action $a\in A$. The weights of the network are updated using the Bellman equation and gradient descent.
@@ 3.1. Imitation learning and unknotting
We used behavioural cloning based on a NN to predict for each crossing the probability that it lies in a minimal unknotting set. If the predicted probability for a crossing $c$ is larger than $0.5$, then we interpret this such that $c$ does lie in a minimal unknotting set. Expert data was obtained from brute-forced minimal unknotting sets of randomly generated knots up to 30 crossings.
The distribution of the percentage of crossings that lie in a minimal unknotting set for diagrams in the dataset is shown in Figure 1. The unknotting numbers of the diagrams range between 1 and 8.
@@
To find out which subset of invariants to use, we set up a Supervised Learning experiment to predict whether a crossing lies in a minimal unknotting sequence. We considered neural networks with hidden layers of size [256, 256], [1024, 1024, 1024], and [2048, 2048, 2048], respectively. Learning was done in 10k steps, with learning rate 0.01 or 0.005, momentum 0.9, and batch size 2048. Accuracy ranged between 84.6% and 88.1%, with baseline 50.1% when predicting the most commonly occurring crossing in unknotting sets (crossing 2 is in a minimal unknotting set for 50.1% of diagrams in the dataset). Note that the indexing of the crossings is arbitrary in the dataset. The highest accuracy was achieved by hidden layers of size [1024,1024,1024] and learning rate 0.01 using all features. When only using the features in (3) on the same architecture, we obtained almost identical, 87.97% accuracy.
@@ 3.2. Reinforcement learning and unknotting
Initially, we observed that the agent frequently revisited the same diagram. If a crossing appeared multiple times in an unknotting sequence, we counted it modulo 2. Furthermore, we experimented with disallowing revisiting the same Alexander or Jones polynomial, which decreased the percentage of unsolved diagrams for 50–100-crossing knots from 0.49% to 0.05% when using both the Alexander and the Jones polynomial. There was also a small increase in performance when disallowing Jones revisits only. When using an agent based on Alexander polynomial features only, disallowing Alexander polynomial revisits substantially improved performance, except when solving connected sums with forced inter-component crossings. When we allowed revisits, the agent sometimes found shorter unknotting sequences (after counting the number of times each crossing change was made modulo 2). The Alexander only policy was 10 times faster than the one using the Jones polynomial. Running the agent 10 times on each diagram substantially increased performance.
@@
We also compared the Jones unknotting agent to various baselines. A naive baseline of randomly switching crossings is fast and simple to implement, but unlikely to find minimal unknotting sequences for diagrams with many crossings. Similarly to [12], we developed a Bayesian Optimisation (BO) based random agent. The agent chooses randomly from the following action categories: Reidemeister I move, reverse Reidemeister I move, Reidemeister II move, reverse Reidemeister II move, Reidemeister III move, and crossing switch. The random agent chooses one of these action categories by randomly sampling according to certain weights. The concrete action is then sampled uniformly from all valid moves from that category, e.g. from all valid crossing indices. Each crossing change carries a reward of $-1$ and the reward is 00 otherwise, which is different from the reward function we used for the RL agent. We optimised the sampling weights using BO using the average reward on a heldout validation set as the fitness metric. We limited the weights to be integers between $1$ and $100$. The best fitness was achieved by weight vectors where weights for simplifying moves (Reidemeister I or II) as well as neutral moves (Reidemeister III) were near $100$, while weights for moves which increase the crossing count (reverse Reidemeister I and II) or which incur a reward penalty (crossing switch) were assigned a weight near $1$. Intuitively, this means the optimisation procedure converged to a solution which always attempts to maximally simplify the knot diagram before making a crossing switch. We then implemented further baselines which follow this insight of first simplifying then switching a crossing. Simplify (random) utilizes SnapPy to first simplify the knot diagram before a random crossing is switched. Simplify (min. crossing) always switches the crossing after which the resulting diagram has the least crossings after simplification. Figure 4 shows how different unknotting strategies compare on a test set of $100,000$ diagrams. The RL agent utilizing the Jones polynomial manages to unknot more knots than the other strategies and is also capable of producing more efficient unknotting sequences.